$800 is invested for 5 years at an annual simple interest rate of 14%
how much interest will be earned?
what is the future value of the investment at the end of the 5 year?

Answer by amarjeeth123(569)

Interest=P*t*r/100 where P is the principal t is the time and r is the rate
Plugging in the values,
Interest=800*5*14/100=$560
Future value of the investment=P+I=800+560=$1360.